OpenStudy (anonymous):

Hi, heres a problem I am trying to do in MATLAB. s = (-1/2*R*C) + a*((1/2*R*C)^2 - (1/L*s))^1/2 . My roblem is how can I solve this is s is on both sides of equation??

OpenStudy (anonymous):

s = (-1/2*R*C) + a*((1/2*R*C)^2 - (1/L*s))^1/2

OpenStudy (anonymous):

and I am given all values except s

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ybarrap):

You have $$ s=-\frac{C R}{2}+a \sqrt{\left(\frac{R C}{2}\right)^2-\frac{s}{L}} $$ Therefore, after rearranging terms, $$ \begin{array}{l} \text{With } ~L\neq 0,~ a\neq 0 \\ s=\frac{-\sqrt{a^2 \left(a^2+C^2 L^2 R^2+2 C L R\right)}-a^2-C L R}{2 L} \\ \end{array} $$

OpenStudy (anonymous):

Im a little confused as to what you did there.......

OpenStudy (anonymous):

I dont think I wrote my equation very cleary.....this was a matlab problem. I think I might open the question again with some slight adjustments..

OpenStudy (anonymous):

and I think it should be s= -1/2RC....

OpenStudy (ybarrap):

$$ s=-\frac{C R}{2}+a \sqrt{\left(\frac{R C}{2}\right)^2-\frac{s}{L}}\\ s+\frac{C R}{2}=a \sqrt{\left(\frac{R C}{2}\right)^2-\frac{s}{L}}\\ (s+\frac{C R}{2})^2=a^2\left ( \cfrac{R C}{2}\right )^2-\frac{s}{L}\\ s^2+CR+\cfrac{(CR)^2}{4}=a^2\left ( \cfrac{R C}{2}\right )^2-\frac{s}{L}\\ s^2+\frac{s}{L}+\cfrac{(CR)^2}{4}-a^2 \cfrac{(R C)^2}{4}=0\\ s^2+\frac{s}{L}+\cfrac{(CR)^2}{4}(1-a^2)=0 $$ Now use quadratic formula.

OpenStudy (ybarrap):

Solve for s: s^2+s/L+1/4 C^2 R^2 (1-a^2) = 0 Solve the quadratic equation by completing the square. Subtract 1/4 (1-a^2) C^2 R^2 from both sides: s^2+s/L = -1/4 (C^2 R^2 (1-a^2)) Take one half of the coefficient of s and square it, then add it to both sides. Add 1/(4 L^2) to both sides: s^2+s/L+1/(4 L^2) = 1/(4 L^2)-1/4 C^2 R^2 (1-a^2) Factor the left hand side. Write the left hand side as a square: (s+1/(2 L))^2 = 1/(4 L^2)-1/4 C^2 R^2 (1-a^2) Eliminate the exponent on the left hand side. Take the square root of both sides: s+1/(2 L) = sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2)) or s+1/(2 L) = -sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2)) Look at the first equation: Solve for s. Subtract 1/(2 L) from both sides: s = sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2))-1/(2 L) or s+1/(2 L) = -sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2)) Simplify powers and products. Combine C^2 and R^2 under the same power: s = sqrt(1/(4 L^2)+1/4 C^2 R^2 (a^2-1))-1/(2 L) or s+1/(2 L) = -sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2)) Look at the second equation: Solve for s. Subtract 1/(2 L) from both sides: s = sqrt(1/(4 L^2)+1/4 C^2 R^2 (a^2-1))-1/(2 L) or s = -1/(2 L)-sqrt(1/(4 L^2)-1/4 C^2 R^2 (1-a^2)) Simplify powers and products. Combine C^2 and R^2 under the same power: Answer: s = sqrt(1/(4 L^2)+1/4 C^2 R^2 (a^2-1))-1/(2 L) or s = -1/(2 L)-sqrt(1/(4 L^2)+1/4 C^2 R^2 (a^2-1))